OpenStudy (anonymous):

What are the coordinates of the vertex for f(x) = 2x2 + 4x + 9? (−2, 4) (−1, 7) (1, 15) (2, 9)

Nnesha (nnesha):

formula to find x coordinate of vertex \[\huge\rm x=\frac{ -b }{ 2a }\]

Nnesha (nnesha):

then substitute all x in f(x) equation for its value to find y-coordinate of vertex

OpenStudy (anonymous):

what do i put for the a and b?

Nnesha (nnesha):

a=leading coefficient b=middle term \[\huge\rm \color{reD}{A}x^2+\color{blue}{B}x+C=0\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

do i leave the negative sign in front of the b even when it's a positive?

Nnesha (nnesha):

yes b is positive but negative is in the formula so that stays there

OpenStudy (anonymous):

so it would be \[x=\frac{ -4 }{ 2(2)}\]?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

so -4/4 and i would just divide it and it would come out as -1 right?

Nnesha (nnesha):

yes that's x-coordinate

OpenStudy (anonymous):

how do i find the y-coordinate now?

Nnesha (nnesha):

now to find y -coordinate substitute x for -1 \[\rm \color{reD}{x=-1}\] \[\huge\rm 2\color{reD}{x}^2+4\color{red}{x}+9=0\]

OpenStudy (anonymous):

so the y-coordinate would be 3? (-1,3)?

Nnesha (nnesha):

mhm

Nnesha (nnesha):

when you take square of negative number you will get positive number

OpenStudy (anonymous):

how do i do that

Nnesha (nnesha):

\[\huge\rm 2\color{reD}{(-1)}^2+4\color{red}{(-1)}+9=0\] now solve (-1)^2 is same as -1 times -1 = ?

OpenStudy (anonymous):

so 1?

Nnesha (nnesha):

yes -1 times -1 = 1 (-1)^2=1 so solve \[\huge\rm 2\color{reD}{(1)}+4\color{red}{(-1)}+9=?\]

OpenStudy (anonymous):

oh so (-1,7)

Nnesha (nnesha):

yep

OpenStudy (anonymous):

could you help me out with another one

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

i'll try

OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (5, −1) and a directrix of y = 1? f(x) = −one fourth (x − 5)2 + 1 f(x) = one fourth (x − 5)2 + 1 f(x) = −one fourth (x − 5)2 f(x) = one fourth (x − 5)2

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

well not sure about this need to review this stuff hmm

Nnesha (nnesha):

sorry

OpenStudy (anonymous):

its fine. i got it Cx thank you tho !

Nnesha (nnesha):

np